Based on data from a​ college, scores on a certain test are normally distributed with a mean of 1519 and a standard deviation of 316.

Find the percentage of scores less than 1298. round to two decimal places as needed
Answer #1

Given,

= 1519 , = 316

We convert this to standard normal as

P(X < x) = P(Z < ( x - ) / )

So,

P(X < 1298) = P(Z < ( 1298 - 1519 ) / 316 )

= P(Z < -0.70)

= 0.2422 (From Z table)

= 24.22%